How would i figure this out,
"Find the remaining lengths of the two sides. AngleB=52 degrees, side a=5 m"
Thank you!

Answer by venugopalramana(3286)

How would i figure this out,
"Find the remaining lengths of the two sides. AngleB=52 degrees, side a=5 m"
Thank you!
A TRIANGLE NEEDS 3 DIMENSIONS TO BE FIXED.HERE WE HAVE ONLY 2 DIMENSIONS.SO THERE WILL BE SEVERAL SUCH TRIANGLES AND HENCE INFINITE ANSWERS.
FOR EXAMPLE
WE ARE GIVEN..ANGLE ABC=52 DEG.....BC=5 M.
CASE 1
TAKE ANGLE BAC ALSO =52...THEN......CA=BC=5 M...ANGLE ACB=180-52-52=76
AB CAN BE FOUND FROM
AB/SIN(76)=5/SIN(52) ....AB=6.16...
CASE 2
TAKE ANGLE BAC=90 DEG....THEN ANGLE ACB=180-90-52=38
5/SIN(90)=BC=CA/SIN(52)=AB=/SIN(38).....
CA=5SIN(52)=3.94
AB=5SIN(38)=3.08
ETC....
YOU CAN HAVE INFINITE ANSWERS